### v1

@@ 4.3 Baseline Systems
Qwen2-Audio (Chu et al., 2024) is a powerful general-purpose audio understanding model capable of performing various audio-related tasks, including the S2TIF task. We also build a cascaded system with Qwen2-Audio and VITS to complete the S2SIF task.
4.4 Main Results
Table 1: ChatGPT scores for S2TIF and S2SIF tasks on the InstructS2S-Eval benchmark, along with the alignment scores between speech and text responses. Here we set $\Omega=+\infty$ for the S2SIF task.
Model S2TIF S2SIF Alignment
Content $\uparrow$ Style $\uparrow$ Content $\uparrow$ Style $\uparrow$ ASR-WER $\downarrow$ ASR-CER $\downarrow$
SpeechGPT 2.59 3.15 1.58 1.81 47.62 37.26
SALMONN (+TTS) 2.57 2.79 2.46 2.84 21.77 20.25
Qwen2-Audio (+TTS) 2.73 2.64 2.32 2.58 55.72 53.09
LLaMA-Omni 3.23 3.81 2.69 3.12 11.61 7.59
Table 1 presents the main results on the InstructS2S-Eval benchmark. First, for the S2TIF task, from the content perspective, LLaMA-Omni shows significant improvement compared to previous models. This is mainly because LLaMA-Omni is developed based on the latest Llama-3.1-8B-Instruct model, leveraging its strong text instruction-following capabilities. From the style perspective, SALMONN and Qwen2-Audio receive lower scores, as they are speech-to-text models. Their output style is not aligned with speech interaction scenarios, often producing formatted content and containing a lot of redundant explanations. In contrast, SpeechGPT, as a speech-to-speech model, achieves a higher style score. Similarly, our LLaMA-Omni attains the highest style score, indicating that after being trained on our InstructS2S-200K dataset, the output style has been well-aligned with speech interaction scenarios. For the S2SIF task, LLaMA-Omni also outperforms previous models in both content and style scores. This further confirms that LLaMA-Omni is capable of effectively addressing user’s instructions with speech in a concise and efficient manner.
Additionally, in terms of alignment between speech and text responses, LLaMA-Omni achieves the lowest ASR-WER and ASR-CER scores. In contrast, SpeechGPT performs poorly in aligning speech and text responses, likely due to its sequential generation of text and speech. The speech-text alignment of cascaded systems, such as SALMONN+TTS and Qwen2-Audio+TTS, is also suboptimal, primarily because the generated text responses may contain characters that cannot be synthesized into speech. This issue is especially evident in Qwen2-Audio, which occasionally outputs Chinese characters, introducing errors in the speech responses. In comparison, LLaMA-Omni achieves the lowest ASR-WER and ASR-CER scores, demonstrating a higher degree of alignment between generated speech and text responses, and further validating the advantage of our approach in simultaneously generating both text and speech responses.
4.5 Trade-off between Speech Quality and Response Latency
Table 2: Latency, speech-text alignment and speech quality under different unit chunk sizes.
Chunk Size $\Omega$ Latency (ms) #Lagging Word ASR-WER $\downarrow$ ASR-CER $\downarrow$ UTMOS $\uparrow$
10 226.13 1.82 10.44 6.94 3.2304
20 256.28 2.83 10.51 6.98 3.4748
40 326.63 4.68 10.99 7.24 3.6688
60 391.96 6.47 11.17 7.27 3.7549
80 467.34 8.29 11.40 7.45 3.7858
100 527.64 10.02 11.48 7.45 3.8242
+$\infty$+ ∞ 1924.62 41.40 11.61 7.59 3.9296
LLaMA-Omni can simultaneously generate both text responses and discrete units corresponding to the speech response. As described in Section 2.6, to further enable streaming waveform generation, when the number of generated discrete units reaches a certain chunk size $\Omega$, the unit chunk is fed into the vocoder to synthesize and play the speech. By adjusting the value of $\Omega$, we can control the system’s latency, where a smaller $\Omega$ corresponds to lower system latency. When $\Omega=+\infty$, it equates to waiting for all units to be generated before synthesizing the speech. At the same time, the value of $\Omega$ also affects the quality of the generated speech. A smaller $\Omega$ means that the speech is divided into more segments for synthesis, which may result in discontinuities between the segments, potentially reducing the overall coherence of the speech.
To better understand the impact of $\Omega$, we explore the system’s latency, the alignment between speech and text responses, and the quality of the generated speech under different $\Omega$ settings. As shown in Table 2, when $\Omega$ is set to 10, the system’s response latency is as low as 226ms, which is even lower than GPT-4o’s average audio latency of 320ms. At this point, the speech response lags by an average of 1.82 words at the start. When $\Omega$ is set to +$\infty$+ ∞, the latency increases to around 2 seconds. For the ASR-WER and ASR-CER metrics, we are surprised to find that as the chunk size increases, the error rates also increase. We believe there may be two reasons for this. On one hand, the vocoder may handle short unit sequences more reliably than long ones, as it is typically trained on shorter sequences. On the other hand, the ASR model we use, Whisper-large-v3, has strong robustness. Even when the speech is somewhat discontinuous with smaller $\Omega$, it has little impact on ASR recognition accuracy. Therefore, we further evaluate the naturalness of the generated speech using the UTMOS metric. It shows that as $\Omega$ increases, the naturalness of the speech improves, since discontinuities in the speech decrease. In summary, we can adjust the value of $\Omega$ based on different scenarios to achieve a trade-off between response latency and speech quality.
4.6 Decoding Time
Table 3: Average decoding time (s) of different models on S2TIF and S2SIF tasks.
Model S2TIF S2SIF
SpeechGPT 4.28 25.60
SALMONN 4.78 /
Qwen2-Audio 8.42 /
LLaMA-Omni 1.49 1.92
Table 3 lists the average decoding times of different models on the S2TIF and S2SIF tasks. For the S2TIF task, SpeechGPT needs to first output the text instruction and then the text response, while SALMONN and Qwen2-Audio tend to produce lengthy responses. In contrast, LLaMA-Omni provides concise answers directly, resulting in significantly lower decoding times, with an average of only 1.49 seconds per instruction. For the S2SIF task, SpeechGPT serially outputs the text and speech responses, resulting in a decoding time approximately 6 times longer than when generating only text responses. In contrast, LLaMA-Omni outputs both text and speech responses simultaneously and employs a non-autoregressive architecture for generating discrete units. As a result, the total generation time only increases by 1.28 times, demonstrating LLaMA-Omni’s advantage in decoding speed.
4.7 Case Study
@@
SALMOON To wrap a present neatly, you can start by laying out a clean and flat surface to work on. Then, place the present on the surface and wrap it in a piece of wrapping paper that is slightly larger than the present. Fold the paper in half and then in half again, so that the top and bottom flaps meet in the middle. Then, fold the left and right flaps in towards the center, so that they meet at the top and bottom of the present. Finally, seal the present by pressing the flaps together and using some tape to secure them.
@@
5 Related Work
Speech/Audio Language Models
With the success of language models in the field of natural language processing (Brown et al., 2020), researchers have begun exploring how to model speech or audio using language models. Early work attempted to train language models on semantic tokens or acoustic tokens of audio, enabling the generation of audio without the need for text (Lakhotia et al., 2021; Nguyen et al., 2023; Borsos et al., 2023). Furthermore, by jointly training speech tokens and text, decoder-only models like VALL-E (Wang et al., 2023b) and VioLA (Wang et al., 2023c) can perform tasks such as speech recognition, speech translation, and speech synthesis. However, the above models are not built upon LLMs. To harness the power of LLMs, many studies explore how to build speech-language models based on LLMs like LLaMA, which can be further divided into two types. The first type, represented by SpeechGPT (Zhang et al., 2023; 2024a) and AudioPaLM (Rubenstein et al., 2023), involves creating native multimodal speech-text models by adding speech tokens to the LLM’s vocabulary and continuing pretraining using speech and text data. However, this approach typically requires a large amount of data and substantial computational resources. The second type typically involves adding a speech encoder before the LLM and finetuning the entire model to equip it with speech understanding capabilities (Shu et al., 2023; Deshmukh et al., 2023), such as speech recognition (Fathullah et al., 2024a; Yu et al., 2024; Ma et al., 2024c; Hono et al., 2024), speech translation (Wu et al., 2023; Wang et al., 2023a; Chen et al., 2024), or other general speech-to-text tasks (Chu et al., 2023; Tang et al., 2024; Chu et al., 2024; Fathullah et al., 2024b; Das et al., 2024; Hu et al., 2024). However, these approaches typically focus only on speech or audio understanding without the ability to generate them. Compared to previous work, LLaMA-Omni equips the LLM with both speech understanding and generation capabilities, enabling it to perform general speech instruction-following tasks. Additionally, LLaMA-Omni has a low training cost, making it convenient for development based on the latest LLMs.
Simultaneous Generation
Streaming generation aims to begin producing output before the entire input is received. This capability is crucial for maintaining synchronization between speakers and listeners in various scenarios, such as streaming speech recognition and simultaneous interpretation. In the case of large language models, having a streaming speech synthesis component can significantly reduce latency between the model and its users. Popular streaming generation methods fall into three main categories: monotonic-attention-based methods (Raffel et al., 2017), CTC-based methods (Graves et al., 2006b), and Transducer-based methods (Graves, 2012). Monotonic-attention-based methods modify the traditional attention-based sequence-to-sequence framework (Bahdanau, 2014) to support streaming generation. These methods rely on an external module to manage the READ/WRITE policy, which can be either fixed (e.g., Wait-k (Ma et al., 2018)) or adaptive (e.g., MMA (Ma et al., 2019), EDAtt (Papi et al., 2022), Seg2Seg (Zhang & Feng, 2024)). CTC-based methods add a blank symbol to the target vocabulary to represent a WAIT action. Streaming inference is achieved by removing adjacent repetitive tokens and blank symbols. To leverage the strengths of attention-based methods, CTC-based approaches often use a chunk-based non-autoregressive architecture (Ma et al., 2023), which has proven effective in simultaneous interpretation and streaming speech synthesis (Zhang et al., 2024b; Ma et al., 2024a). Transducer-based methods are designed to bridge the gap between the non-autoregressive nature of CTC-based methods and the autoregressive dependency between target tokens. These approaches introduce an additional predictor to capture token dependencies, and their variants have shown strong performance in simultaneous interpretation (Liu et al., 2021; Tang et al., 2023). More recently, researchers have begun adopting decoder-only large language models for streaming generation tasks (Seide et al., 2024; Guo et al., 2024), and extending them to an interruptible duplex model (Ma et al., 2024b).
6 Conclusion
In this paper, we propose an innovative model architecture, LLaMA-Omni, which enables low-latency and high-quality speech interaction with LLMs. LLaMA-Omni is built upon the latest Llama-3.1-8B-Instruct model, with the addition of a speech encoder for speech understanding and a streaming speech decoder that can generate both text and speech responses simultaneously. To align the model with speech interaction scenarios, we construct a speech instruction dataset InstructionS2S-200K, which contains 200K speech instructions along with the speech responses. Experimental results show that, compared to previous speech-language models, LLaMA-Omni delivers superior responses in both content and style, with a response latency as low as 226ms. Moreover, training LLaMA-Omni requires less than 3 days on 4 GPUs, enabling rapid development of speech interaction models based on the latest LLMs. In the future, we plan to explore enhancing the expressiveness of generated speech responses and improving real-time interaction capabilities.
Appendix A Prompt
Prompt for ChatGPT Scoring (Model: GPT-4o) I need your help to evaluate the performance of several models in the speech interaction scenario. The models will receive a speech input from the user, which they need to understand and respond to with a speech output. Your task is to rate the model’s responses based on the provided user input transcription [Instruction] and the model’s output transcription [Response]. Please evaluate the response from two perspectives: content and style, and provide a score for each on a scale of 1 to 5. Content (1-5 points): 1 point: The response is largely irrelevant, incorrect, or fails to address the user’s query. It may be off-topic or provide incorrect information. 2 points: The response is somewhat relevant but lacks accuracy or completeness. It may only partially answer the user’s question or include extraneous information. 3 points: The response is relevant and mostly accurate, but it may lack conciseness or include unnecessary details that don’t contribute to the main point. 4 points: The response is relevant, accurate, and concise, providing a clear answer to the user’s question without unnecessary elaboration. 5 points: The response is exceptionally relevant, accurate, and to the point. It directly addresses the user’s query in a highly effective and efficient manner, providing exactly the information needed. Style (1-5 points): 1 point: The response is poorly suited for speech interaction, possibly including structured elements like lists or being overly complex, disjointed, or difficult to understand. 2 points: The response is somewhat suitable but may be too long, too short, or awkwardly phrased, making it less effective in a speech interaction context. 3 points: The response is generally suitable for speech interaction, but it may have minor issues with length, clarity, or fluency that detract slightly from the overall effectiveness. 4 points: The response is well-suited for speech interaction, with appropriate length, clear language, and a natural flow. It is easy to understand when spoken aloud. 5 points: The response is perfectly suited for speech interaction. It is the ideal length, highly clear, and flows naturally, making it easy to follow and understand when spoken. Below are the transcription of user’s instruction and models’ response: ### [Instruction]: {instruction} ### [Response]: {response} After evaluating, please output the scores in JSON format: {“content”: content score, “style”: style score}. You don’t need to provide any explanations.

### v2

@@ 4.3 Baseline Systems
Qwen2-Audio (Chu et al., 2024) is a powerful general-purpose audio understanding model capable of performing various audio-related tasks, including the S2TIF task. We also build a cascaded system with Qwen2-Audio and Orca to complete the S2SIF task.
When using Orca for streaming speech synthesis, we need to set a word chunk size $\Theta$, which means that speech synthesis is triggered every time $\Theta$ new words arrive. In our experiments, we varies $\Theta$ within the range of [1, 3, 5, 7, 9] to control the response latency of cascaded systems.
4.4 Results in the Offline Scenario
Table 2 presents the results on the InstructS2S-Eval benchmark in the offline scenario. For the S2TIF task, LLaMA-Omni achieves the highest ChatGPT Score, significantly outperforming the baseline systems. We attribute this to two key factors. First, our model is built upon the latest Llama-3.1-8B-Instruct model, leveraging its strong text instruction-following capabilities. Second, our InstructS2S-200K dataset effectively aligns the model with speech interaction scenarios, ensuring that its responses are both high-quality and well-suited to speech contexts. In comparison, SALMONN and Qwen2-Audio, as speech-to-text models, have not been aligned with speech interaction scenarios. As a result, their responses often include formatted content and a lot of redundant explanations, making them less suitable for such contexts.
Table 2: Results on the InstructS2S-Eval benchmark in the offline scenario. $\Delta$ refers to the difference in ChatGPT Score between the S2SIF and S2TIF tasks.
Model ChatGPT Score ASR-WER $\downarrow$ UTMOS $\uparrow$
S2TIF S2SIF $\Delta$
SpeechGPT 2.98 2.19 0.79 45.00 3.8958
SALMONN + Orca 3.44 3.40 0.04 3.78 3.8286
Qwen2-Audio + Orca 3.47 3.38 0.09 6.77 3.6119
LLaMA-Omni 3.99 3.47 0.52 10.82 3.9296
(a) ChatGPT Score against latency.
(b) ASR-WER against latency.
(c) UTMOS against latency.
(d) WPS against latency.
Figure 3: Results on the InstructS2S-Eval benchmark in the streaming scenario. We report the ChatGPT Score, ASR-WER, UTMOS, and WPS under different latency conditions.
For the S2SIF task, LLaMA-Omni still achieves the highest ChatGPT score among all models. We observe that scores for the S2SIF task are generally lower than those for the S2TIF task. This decline is primarily due to errors introduced during the speech synthesis process and the reliance on an ASR model for evaluation. The two cascaded baseline systems, which use industrial TTS models for speech synthesis, exhibit the smallest decline in scores. For the end-to-end models, LLaMA-Omni demonstrates a smaller drop in score compared to SpeechGPT, indicating that LLaMA-Omni has stronger speech generation capabilities. This can be further verified by the ASR-WER metric: the cascaded systems achieve the lowest ASR-WER, while LLaMA-Omni’s ASR-WER is slightly higher but still acceptable. In contrast, SpeechGPT’s ASR-WER is significantly higher, suggesting poor alignment between its speech and text responses. Although the alignment between LLaMA-Omni’s speech and text responses is slightly lower than that of the cascaded systems, this is primarily due to the fact that LLaMA-Omni’s speech decoder is trained on only approximately 1K hours of data, which is far less than the industrial TTS model. We believe that with more training data, LLaMA-Omni’s performance could be further improved. Finally, the UTMOS metric shows that LLaMA-Omni generates satisfactory speech quality, slightly surpassing other baseline models.
4.5 Results in the Streaming Scenario
In the streaming scenario, the model generates speech responses simultaneously while generating text responses. For SpeechGPT and LLaMA-Omni, the response latency is controlled by adjusting the unit chunk size $\Omega$, whereas for cascaded systems, it is managed by adjusting the word chunk size $\Theta$. In Figure 3, we present the ChatGPT Score, ASR-WER, UTMOS and WPS results of all models under different latency conditions. Firstly, we examine the results of LLaMA-Omni under different latency conditions and observe that LLaMA-Omni can achieve a minimum latency of 236ms (with $\Omega=10$), which is even lower than GPT-4o’s average audio latency of 320ms. As latency increases ($\Omega$ grows larger), we notice a slight decrease in the ChatGPT Score, while ASR-WER shows a minor improvement. We believe this is mainly because the vocoder may handle shorter unit sequences more reliably than longer ones, as it is typically trained on shorter sequences. However, when $\Omega$ is smaller, the speech is divided into more segments for synthesis, which increases discontinuities in the speech. This leads to a decline in speech quality, and consequently, a decrease in the UTMOS score is observed. In addition, the speech generated by LLaMA-Omni maintains almost consistent speech rate across different latency conditions. In summary, LLaMA-Omni achieves relatively stable performance under varying latency conditions. At a latency of 563ms ($\Omega=100$), the results of all metrics are relatively close to those in the offline scenario.
Next, we compare LLaMA-Omni with other baseline systems. For SpeechGPT, its decoding latency is significantly higher than other systems ($>$4500ms) due to the sequential generation of text instructions and text responses before generating units. Additionally, the quality of its generated responses is relatively poor. For cascade systems based on streaming TTS, we observe that both Qwen2-Audio + Orca and SALMONN + Orca can achieve latencies below 300ms. However, they exhibit two notable issues. Firstly, in the streaming scenario, their ChatGPT Score and ASR-WER show significant gaps compared to the offline scenario, indicating that streaming speech synthesis at the word level tends to introducing additional errors. Secondly, as shown in Figure 3(d), when the latency is low, the overall speech rate of cascaded systems drops sharply. This is mainly because the pauses between words become more frequent, resulting in reduced naturalness and coherence of the speech. In contrast, LLaMA-Omni’s streaming unit generation is performed within the end-to-end model, requiring only a cascaded vocoder to complete the streaming unit-to-waveform conversion. As a result, the overall prosody and rhythm of the speech remain almost unchanged across different latency levels. This demonstrates the advantage of end-to-end models over cascade systems based on streaming TTS. This can be more intuitively experienced by listening to the provided audio samples. We provide the numerical results of all models in the streaming scenario in Appendix C.
4.6 Human Evaluation
(a) Helpfulness.
(b) Naturalness.
Figure 4: Results of human evaluation from the perspectives of helpfulness and naturalness.
To better understand human preferences for model responses in real-time speech interaction scenarios, we further conduct a human evaluation. Specifically, we perform side-by-side comparisons between LLaMA-Omni ($\Omega=40$, latency=347ms) and two cascaded systems: SALMONN+Orca ($\Theta=3$, latency=352ms) and Qwen2-Audio+Orca ($\Theta=3$, latency=420ms). For each comparison, we randomly select 20 speech instructions and collect the speech responses generated by the two models. We then invite 5 participants to evaluate all samples from two key perspectives: helpfulness and naturalness. Helpfulness evaluates whether the model follows the instructions and provides appropriate responses, while naturalness evaluates the fluency and natural quality of the generated speech. Participants compare each pair of samples rate them as win, tie, or lose for both aspects. Finally, each pairwise model comparison results in a total of 100 votes. The results, shown in Figure 5, demonstrate that LLaMA-Omni achieves a higher win rate compared to the cascade systems in both helpfulness and naturalness, confirming that LLaMA-Omni generates responses that better align with human preferences.
5 Related Work
Speech/Audio Language Models
With the success of language models in the field of natural language processing (Brown et al., 2020), researchers have begun exploring how to model speech or audio using language models. Early work attempted to train language models on semantic tokens or acoustic tokens of audio, enabling the generation of audio without the need for text (Lakhotia et al., 2021; Nguyen et al., 2023; Borsos et al., 2023). Furthermore, by jointly training speech tokens and text, decoder-only models like VALL-E (Wang et al., 2023b) and VioLA (Wang et al., 2023c) can perform tasks such as speech recognition, speech translation, and speech synthesis. However, the above models are not built upon LLMs. To harness the power of LLMs, many studies explore how to build speech-language models based on LLMs like LLaMA, which can be further divided into two types. The first type, represented by SpeechGPT (Zhang et al., 2023; 2024a), AudioPaLM (Rubenstein et al., 2023), and AnyGPT (Zhan et al., 2024), involves creating native multimodal speech-text models by adding speech tokens to the LLM’s vocabulary and continuing pretraining using speech and text data. However, this approach typically requires a large amount of data and substantial computational resources. The second type typically involves adding a speech encoder before the LLM and finetuning the entire model to equip it with speech understanding capabilities (Shu et al., 2023; Deshmukh et al., 2023), such as speech recognition (Fathullah et al., 2024a; Yu et al., 2024; Ma et al., 2024b; Hono et al., 2024), speech translation (Wu et al., 2023; Wang et al., 2023a; Chen et al., 2024), or other general speech-to-text tasks (Chu et al., 2023; Tang et al., 2024; Chu et al., 2024; Fathullah et al., 2024b; Das et al., 2024; Hu et al., 2024). However, these approaches typically focus only on speech or audio understanding without the ability to generate them. Compared to previous work, LLaMA-Omni equips the LLM with both speech understanding and generation capabilities, enabling it to perform general speech instruction-following tasks. Recently, some contemporary works such as Mini-Omni (Xie & Wu, 2024) and Moshi (Défossez et al., 2024) have also focused on speech interactions with LLMs, aiming to improve the quality of speech response by simultaneously generating both text and speech. Compared to them, our advantages include: (1) LLaMA-Omni is built upon the latest LLM Llama-3.1-8B-Instruct model, which has stronger reasoning capabilities; (2) We utilize CTC to adaptively learn the alignment between speech and text responses, eliminating the need to pre-align speech and text during training; (3) Our training only uses 200K data samples, which is several orders of magnitude less than theirs, and the training requires only 4 GPUs for 3 days, making the training cost significantly lower.
Simultaneous Generation
Streaming generation aims to begin producing output before the entire input is received. This capability is crucial for maintaining synchronization between speakers and listeners in various scenarios, such as streaming speech recognition and simultaneous interpretation. In the case of large language models, having a streaming speech synthesis component can significantly reduce latency between the model and its users. Popular streaming generation methods fall into three main categories: monotonic-attention-based methods (Raffel et al., 2017), CTC-based methods (Graves et al., 2006b), and Transducer-based methods (Graves, 2012). Monotonic-attention-based methods modify the traditional attention-based sequence-to-sequence framework (Bahdanau, 2014) to support streaming generation. These methods rely on an external module to manage the READ/WRITE policy, which can be either fixed (e.g., Wait-k (Ma et al., 2018)) or adaptive (e.g., MMA (Ma et al., 2019), EDAtt (Papi et al., 2022), Seg2Seg (Zhang & Feng, 2024)). CTC-based methods add a blank symbol to the target vocabulary to represent a WAIT action. Streaming inference is achieved by removing adjacent repetitive tokens and blank symbols, which has proven effective in simultaneous interpretation and streaming speech synthesis (Ma et al., 2023; Zhang et al., 2024b; Ma et al., 2024a). Transducer-based methods are designed to bridge the gap between the non-autoregressive nature of CTC-based methods and the autoregressive dependency between target tokens. These approaches introduce an additional predictor to capture token dependencies, and their variants have also shown strong performance in simultaneous interpretation (Liu et al., 2021; Tang et al., 2023) and streaming speech synthesis (Chen et al., 2021). For the streaming TTS task, many studies employ straightforward lookahead strategies built upon models like Tacotron 2 (Shen et al., 2018), such as waiting for a few future words (Ma et al., 2020; Stephenson et al., 2020) or leveraging a language model to predict several future words (Saeki et al., 2021a; b; Liu et al., 2022), enabling improved streaming speech synthesis. Dekel et al. (2024) achieves streaming speech synthesis simultaneously with the output stream of LLMs by employing two cascaded streamable models that sequentially generate phonemes and speech in real-time.
6 Conclusion
In this paper, we propose an innovative model architecture, LLaMA-Omni, which enables low-latency and high-quality speech interaction with LLMs. LLaMA-Omni is built upon the latest Llama-3.1-8B-Instruct model, with the addition of a speech encoder for speech understanding and a streaming speech decoder that can generate both text and speech responses simultaneously. To align the model with speech interaction scenarios, we construct a speech instruction dataset InstructionS2S-200K, which contains 200K speech instructions along with the speech responses. Experimental results show that, compared to previous speech-language models, LLaMA-Omni delivers superior responses in both content and style, with a response latency as low as 236ms. Moreover, training LLaMA-Omni requires less than 3 days on 4 GPUs, enabling rapid development of speech interaction models based on the latest LLMs. In the future, we plan to explore enhancing the expressiveness of generated speech responses and improving real-time interaction capabilities.
Acknowledgments
We thank all the anonymous reviewers for their insightful and valuable comments. This paper is supported by National Natural Science Foundation of China (Grant No.62376260).
Appendix A Prompt
Prompt template of LLaMA-Omni <|begin_of_text|><|start_header_id|>system<|end_header_id|> You are a helpful language and speech assistant. You are able to understand the speech content that the user provides, and assist the user with a variety of tasks using natural language. <|eot_id|> <|start_header_id|>user<|end_header_id|> <speech> Please answer the questions in the user’s input speech. <|eot_id|> <|start_header_id|>assistant<|end_header_id|>
Prompt for ChatGPT Scoring (Model: GPT-4o) I need your help to evaluate the performance of several models in a speech interaction scenario. The models receive the user’s speech input and respond with speech output. For evaluation purposes, both the user’s speech input and the model’s speech response have been transcribed into text using Automatic Speech Recognition (ASR). Your task is to rate the model’s responses based on the provided user input transcription [Instruction] and the model’s output transcription [Response]. Please consider factors such as helpfulness, relevance, fluency, and suitability for speech interaction in your evaluation, and provide a single score on a scale from 1 to 5. Below are the transcription of user’s instruction and models’ response: ### [Instruction]: {instruction} ### [Response]: {response} After evaluating, please output the scores in JSON format: {score: …}. You don’t need to provide any explanations.
Appendix B Case Study
@@
SALMONN To wrap a present neatly, you can start by laying out a clean and flat surface to work on. Then, place the present on the surface and wrap it in a piece of wrapping paper that is slightly larger than the present. Fold the paper in half and then in half again, so that the top and bottom flaps meet in the middle. Then, fold the left and right flaps in towards the center, so that they meet at the top and bottom of the present. Finally, seal the present by pressing the flaps together and using some tape to secure them.
@@
Appendix C Numerical Results in the Streaming Scenario
We present the numerical results of all models in streaming scenarios in Tables 4, 5, 6, and 7.
Table 4: Numerical results of LLaMA-Omni in the streaming scenario.
$\mathbf{\Omega}$ Latency (ms) ChatGPT Score ASR-WER UTMOS WPS
LLM Vocoder Total
10 206.03 30.15 236.18 3.54 9.84 3.2304 2.76
20 236.18 45.23 281.41 3.56 9.91 3.4748 2.75
40 301.51 45.23 346.73 3.52 10.37 3.6688 2.74
60 361.81 50.25 412.06 3.52 10.47 3.7549 2.74
80 432.16 55.28 487.44 3.50 10.70 3.7858 2.73
100 497.49 65.33 562.81 3.49 10.71 3.8242 2.74
Offline 1542.71 211.06 1753.77 3.47 10.82 3.9296 2.73
Table 5: Numerical results of SpeechGPT in the streaming scenario.
$\mathbf{\Omega}$ Latency (ms) ChatGPT Score ASR-WER UTMOS WPS
LLM Vocoder Total
10 4899.50 30.15 4929.65 2.16 44.85 2.7099 1.95
20 5005.03 30.15 5035.18 2.22 42.03 3.1920 1.86
40 5547.74 40.20 5587.94 2.17 43.88 3.5106 1.89
60 6005.03 45.23 6050.25 2.16 43.35 3.6293 1.84
80 6160.80 55.28 6216.08 2.21 43.98 3.6970 1.87
100 8301.51 65.33 8366.83 2.20 43.74 3.7397 1.86
Offline 17919.60 170.85 18090.45 2.19 45.00 3.8958 1.85
Table 6: Numerical results of SALMONN + Orca in the streaming scenario.
$\mathbf{\Theta}$ Latency (ms) ChatGPT Score ASR-WER UTMOS WPS
LLM TTS Total
1 212.45 19.71 232.16 3.28 6.64 3.0947 1.86
3 316.59 35.08 351.67 3.09 8.65 3.7338 2.91
5 428.79 32.08 460.87 3.04 9.23 3.7750 3.18
7 536.47 45.90 582.37 3.06 8.49 3.7972 3.20
9 659.57 69.38 728.95 3.23 7.07 3.8060 3.22
Offline 4274.48 1049.59 5324.07 3.40 3.78 3.8286 3.31
Table 7: Numerical results of Qwen2-Audio + Orca in the streaming scenario.
$\mathbf{\Theta}$ Latency (ms) ChatGPT Score ASR-WER UTMOS WPS
LLM TTS Total
1 289.46 19.15 308.61 2.79 25.25 2.8597 1.90
3 381.66 38.19 419.85 2.95 13.30 3.5529 2.91
5 470.55 38.64 509.19 2.93 13.00 3.5865 3.08
7 568.22 52.95 621.17 3.07 10.59 3.5739 3.11
9 675.55 81.02 756.57 3.14 9.81 3.6016 3.13
Offline 7062.93 2361.49 9424.42 3.38 6.77 3.6119 3.30